# Segment Anything Meets Point Tracking

### v1

@@ 4.4.1 Video Object Segmentation
Qualitative Analysis.
Visualizations of successful video segmentation on DAVIS $2017$ for SAM-PT and SAM-PT-reinit can be seen in Fig. 6(a) and Fig. 6(b) respectively. Notably, Fig. 8 presents successful video segmentation on unseen web videos – clips from the “Avatar: The Last Airbender” anime-influenced animated television series, demonstrating the zero-shot capabilities of our method.
Limitations and Challenges.
Despite the competitive zero-shot performance, certain limitations persist, primarily due to the limitations of our point tracker in handling occlusion, small objects, motion blur, and re-identification. In such scenarios, the point tracker’s errors propagate into future video frames. Fig. 6(c) illustrates these problematic instances on DAVIS $2017$, while Fig. 9 presents additional cases on “Avatar: The Last Airbender” clips. Although using point re-initialization and negative points somewhat alleviates the failures of the point tracker, they still prevent the performance from being on par with methods trained on video data.
Table 4: Quantitative results in semi-supervised VOS on the validation subset of YouTube-VOS $2018$. Metrics are reported separately for “seen” and “unseen” classes, with $\mathcal{G}$ being the overall average score over the metrics. Note that SegGPT and SAM-PT adopt completely different training data. ♠: our reproduced result using the official code of [45].
YouTube-VOS $\mathbf{2018}$ Validation [38]
Method $\mathcal{G}$ $\mathcal{J}_{s}$ $\mathcal{F}_{s}$ $\mathcal{J}_{u}$ $\mathcal{F}_{u}$
(a) trained on video segmentation data
AGAME [13] $66.0$ $66.9$ $61.2$
AGSS [20] $71.3$ $71.3$ $65.5$ $75.2$ $73.1$
STM [25] $79.4$ $79.7$ $84.2$ $72.8$ $80.9$
AFB-URR [18] $79.6$ $78.8$ $83.1$ $74.1$ $82.6$
RDE [17] $83.3$ $81.9$ $86.3$ $78.0$ $86.9$
SWEM [21] $82.8$ $82.4$ $86.9$ $77.1$ $85.0$
XMem [4] $86.1$ $85.1$ $89.8$ $80.3$ $89.2$
SwinB-DeAOT-L [41] $86.2$ $85.6$ $90.6$ $80.0$ $88.4$
@@
Painter [35] $24.1$ $27.6$ $35.8$ $14.3$ $18.7$
SegGPT [36] $\mathbf{74.7}$ $\mathbf{75.1}$ $\mathbf{80.2}$ $\mathbf{67.4}$ $\mathbf{75.9}$
PerSAM-F ♠ [45] $54.4$ $53.9$ $56.4$ $50.7$ $56.6$
SAM-PT (ours) $67.0\pm 0.3$ $68.6\pm 0.2$ $71.2\pm 0.1$ $61.0\pm 0.5$ $67.4\pm 0.4$
SAM-PT-reinit (ours) ${67.5}\pm 0.2$ ${69.0}\pm 0.4$ ${69.9}\pm 0.3$ ${63.2}\pm 0.4$ ${67.8}\pm 0.5$
Table 5: Quantitative results in semi-supervised VOS on the validation subset of MOSE $2023$ [8]. ♠: our reproduced result using the official code of [45].
MOSE $\mathbf{2023}$ Validation [8]
@@
RDE [17] $48.8$ $44.6$ $52.9$
SWEM [21] $50.9$ $46.8$ $54.9$
XMem [4] $57.6$ $53.3$ $62.0$
DeAOT [41] $59.4$ $55.1$ $63.8$
@@
Painter [35] $14.5$ $10.4$ $18.5$
SegGPT [36] $\mathbf{45.1}$ $\mathbf{42.2}$ $\mathbf{48.0}$
PerSAM-F ♠ [45] $23.3$ $19.8$ $26.8$
SAM-PT (ours) $38.5\pm 0.2$ $34.9\pm 0.3$ $42.1\pm 0.2$
SAM-PT-reinit (ours) ${41.0}\pm 0.5$ ${38.5}\pm 0.5$ ${43.5}\pm 0.5$
4.4.2 Video Instance Segmentation
Results and Analysis.
Given the same mask proposals, SAM-PT outperforms TAM [40] significantly even though SAM-PT was not trained on any video segmentation data. TAM is a concurrent approach combining SAM and XMem [4], where XMem was pre-trained on BL30K [5] and trained on DAVIS and YouTube-VOS, but not on UVO. On the other hand, SAM-PT combines SAM with the PIPS point tracking method, both of which have not been trained on video segmentation tasks.
5 Conclusion
We present SAM-PT, an innovative solution that extends SAM’s segmentation ability from static images to dynamic videos. Integrated with long-term point trackers, our approach demonstrates strong performance across several benchmarks including DAVIS, YouTube-VOS, MOSE, and UVO. While our method has limitations such as difficulty handling occlusions, small objects, and motion blur, and inconsistencies in mask predictions, it contributes a simple and effective new point-based perspective to video object segmentation research. By illustrating a promising way to extend foundational models like SAM into the video domain, our research provides a potential pathway for advancements in diverse applications from autonomous driving to video labeling. Furthermore, the future incorporation of more advanced point trackers can enhance the performance of SAM-PT.
@@ 5 Conclusion
Method Propagation $\mathbf{AR100}$ $\mathbf{ARs}$ $\mathbf{ARm}$ $\mathbf{ARl}$ $\mathbf{AP}$
Mask2Former VIS [44] N/A $35.4$- -- -- $27.3$
ROVIS [44] N/A $41.2$- -- -- $32.7$
(b) trained on video segmentation data
TAM [40] Feature Matching $24.1$ $21.1$ $32.9$ $31.1$ $1.7$
(c) not trained on video segmentation data (zero-shot)
SAM-PT (ours) Points Prompting $28.8\pm 0.1$ $23.3\pm 0.1$ $40.8\pm 0.2$ $48.3\pm 0.6$ $\mathbf{6.7}\pm 0.2$
SAM-PT-reinit (ours) Points Prompting $\mathbf{30.8}$ $\mathbf{25.1}$ $\mathbf{44.1}$ $\mathbf{49.2}$ $6.5$
Appendix A Point Tracking Reinitialization

### v2

@@ 4.4 Video Object Segmentation
Qualitative Analysis. Our visualizations in Fig. 4(a) and Fig. 4(b) demonstrate successful segmentation on the DAVIS $2017$ dataset and underscore our method’s ability to perform zero-shot video segmentation on unseen content, such as clips from the anime-influenced series “Avatar: The Last Airbender” in Fig. 6. These examples highlight the versatility and adaptability of SAM-PT.
Limitations and Challenges. Our method excels in zero-shot video object segmentation but faces challenges with point tracker reliability in complex scenarios, such as occlusions and fast-moving objects, as shown in Fig. 4(c). While point reinitialization and negative point strategies offer some improvement, interactive use significantly bridges the performance gap with trained methods. This interactive potential is successfully demonstrated in Sec. 4.6, where user intervention enhances segmentation accuracy.
4.5 Video Instance Segmentation
Given the same mask proposals, SAM-PT outperforms TAM [46] significantly, as shown in Tab. 7, even though SAM-PT was not trained on any video segmentation data. TAM is a concurrent approach combining SAM and XMem [7], where XMem was pre-trained on BL30K [9] and trained on DAVIS and YouTube-VOS, but not on UVO. On the other hand, SAM-PT combines SAM with the PIPS or CoTracker point tracking method, both of which have not been trained on video segmentation tasks.
4.6 Interactive Point-Based Video Segmentation
Building upon our method’s strengths observed in standard benchmarks, this study evaluates the responsiveness of SAM-PT to human input, aiming to understand its practical utility for interactive video annotation tasks. We benchmarked SAM-PT’s interactive performance against a baseline SAM approach that does not utilize point tracking. The results, visualized in Fig. 7, show that SAM-PT significantly outperforms the baseline, particularly when employing the offline checkpoint strategy. Although this offline method initially starts at a lower performance due to setup costs, it quickly exceeds the online method’s performance by benefiting from a cumulative optimization approach.
Figure 7: Interactive segmentation performance on the DAVIS 2017 [30] validation set. SAM-PT’s offline strategy notably outperforms the baseline SAM, demonstrating efficient video annotation with minimal user intervention.
These results suggest that SAM-PT substantially reduces the effort required for high-quality video annotation, bringing its performance closer to fully-supervised methods and highlighting its practical utility.
Table 6: This table shows the semi-supervised VOS performance on BDD100K’s validation set, comparing our HQ-SAM-PT method using CoTracker as the point tracker (without reinitialization) against the zero-shot SegGPT and the fully-supervised XMem. Performance metrics include the average $\mathcal{J\&F}$ measure for object visibility durations categorized as short (1-5 frames), medium (6-30 frames), and long (31+ frames). Our approach demonstrates superior results over SegGPT for non-transient objects and over XMem across all visibility durations except for long-term object tracking.
BDD100K VOS Validation [50]
[…]
(a) trained on video segmentation data, but not on BDD100K
XMem [7] $76.6$ $74.5$ $78.7$ $79.3$ $78.6$ $\mathbf{63.7}$
@@ 4.6 Interactive Point-Based Video Segmentation
SegGPT [42] $\mathbf{81.5}$ $\mathbf{81.2}$ $\mathbf{81.8}$ $\mathbf{96.1}$ $\underline{78.6}$ $52.0$
HQ-SAM-PT (ours) $\underline{81.0}$ $\underline{80.1}$ $\mathbf{81.8}$ $\underline{91.8}$ $\mathbf{79.9}$ $\underline{55.8}$
[…]
Method Tracker Reinit $\mathbf{AR100}$ $\mathbf{ARs}$ $\mathbf{ARm}$ $\mathbf{ARl}$ $\mathbf{AP}$
Mask2Former VIS [51] - - $35.4$- -- -- $27.3$
ROVIS [51] - - $41.2$- -- -- $32.7$
(b) trained on video segmentation data
TAM [46] - - $24.1$ $21.1$ $32.9$ $31.1$ $1.7$
(c) not trained on video segmentation data (zero-shot)
SAM-PT (ours) PIPS [16] $28.8$ $23.3$ $40.8$ $48.3$ $\mathbf{6.7}$
PIPS [16] $\mathbf{30.8}$ $\mathbf{25.1}$ $\mathbf{44.1}$ $\mathbf{49.2}$ $6.5$
CoTracker [19] $29.5$ $25.3$ $39.0$ $44.1$ $5.8$
CoTracker [19] $29.8$ $25.1$ $40.6$ $45.6$ $6.2$
5 Conclusion
SAM-PT introduces a point-centric approach for interactive video segmentation by combining the generalization capabilities of the Segment Anything Model (SAM) and long-term point tracking. Our work fills in the gap that the point-centric approach is scantly explored in the literature. In our experiments, SAM-PT achieves strong performance across video segmentation tasks including semi-supervised, open-world, and fully interactive video segmentation. While our method has limitations such as difficulty handling occlusions, small objects, motion blur, and inconsistencies in mask predictions, it contributes a new perspective to video object segmentation research.
Supplementary Material
In this supplementary, we first report experimental results on additional datasets and subsets (Appendices A, B and C). Then we extend and detail our ablation and report more qualitative results (Appendices D, E, F, G and H). Lastly, we detail on our evaluation protocols (Appendices I, J, K and L).
Appendix A MOSE 2023
MOSE $2023$ [13] is a recently introduced dataset that focuses on multi-object segmentation and tracking in complex scenes, replete with challenges such as occlusions, transient visibility of objects, extensive occlusion, etc. Our results on the validation subset suggest that SAM-PT achieves performance competitive with SegGPT [42], as shown in Tab. 8.
Table 8: Quantitative results on the MOSE $2023$ validation set for semi-supervised VOS. Our method achieves performance comparable to the state-of-the-art zero-shot learning method. Note that SegGPT and SAM-PT adopt completely different training data.
Method Tracker Reinit MOSE $\mathbf{2023}$ Validation [13]
@@ Appendix A MOSE 2023
RDE [23] - - $48.8$ $44.6$ $52.9$
XMem [7] - - $57.6$ $53.3$ $62.0$
DeAOT [48] - - $59.4$ $55.1$ $63.8$
DEVA [10] - - $66.5$ $62.3$ $70.8$
@@
Painter [41] - - $14.5$ $10.4$ $18.5$
SegGPT [42] - - $\mathbf{45.1}$ $\mathbf{42.2}$ $\mathbf{48.0}$
SAM-PT (ours) PIPS [16] $38.5{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.2}}$ $34.9{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.3}}$ $42.1{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.2}}$
PIPS [16] ${41.0}{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.5}}$ ${38.5}{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.5}}$ ${43.5}{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.5}}$
CoTracker [19] $41.8{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.2}}$ $38.3{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.2}}$ $45.2{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.3}}$
CoTracker [19] $40.1{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.5}}$ $36.0{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.5}}$ $44.1{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.4}}$
TAPIR [15] $\underline{42.9}{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.2}}$ $38.3{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.2}}$ $\underline{47.6}{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.1}}$
HQ-SAM-PT (ours) CoTracker [19] $42.4$±${\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.3}}$ $\underline{39.0}$±${\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.3}}$ $45.8$±${\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.3}}$
TAPIR [15] $42.1{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.1}}$ $37.6{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.1}}$ $46.7{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.1}}$
Appendix B DAVIS 2016
DAVIS $2016$ offers a single-object VOS benchmark across $20$ diverse sequences. We report results on the DAVIS $2016$ validation subset in Tab. 9, in which our method achieves $84.3$ points, surpassing SegGPT’s $82.3$ points.
Table 9: Quantitative results on the DAVIS $2016$ validation set for semi-supervised VOS. Our method achieves higher performance compared to SegGPT, both of which are zero-shot methods.
DAVIS 2016 Validation [30]
@@ Appendix B DAVIS 2016
SegGPT [42] - - $82.3$ $81.8$ $82.8$
SAM-PT PIPS [16] 83.183.1{83.1}83.1±$\pm 1.5$ 83.083.0{83.0}83.0±$\pm 0.8$ 83.083.0{83.0}83.0±$\pm 1.1$
PIPS [16] $80.2$±$\pm 0.6$ $80.3$±$\pm 0.6$ $80.0$±$\pm 0.6$
CoTracker [19] $83.1\pm 0.6$ $83.2\pm 0.7$ $82.9\pm 0.6$
CoTracker [19] $82.6\pm 0.8$ $83.0\pm 1.0$ $82.2\pm 0.9$
HQ-SAM-PT CoTracker [19] $\mathbf{84.3}\pm 0.9$ $\mathbf{84.9}\pm 1.0$ $\mathbf{83.7}\pm 0.9$
Appendix C DAVIS 2017 Test-dev Subset
DAVIS $2017$ is a multi-object extension of its $2016$ version. The video scenarios within this dataset are small but diverse. In addition to the results on the validation subset in the main manuscript, we report the performance of SAM-PT on the DAVIS $2017$ test-dev subset in Tab. 10.
Table 10: Quantitative results on the DAVIS 2017 test-dev subset for semi-supervised VOS.
DAVIS 2017 Test-dev [30]
[…]
SAM-PT PIPS [16] 62.762.7{62.7}62.7±$\pm 0.5$ 59.459.4{59.4}59.4±$\pm 0.6$ 66.166.1{66.1}66.1±$\pm 0.4$
PIPS [16] $61.5$±$\pm 1.1$ $59.3$±$\pm 1.0$ $63.8$±$\pm 1.2$
CoTracker [19] $65.7$±$\pm 0.7$ $62.8$±$\pm 0.7$ $68.5$±$\pm 0.7$
CoTracker [19] $62.0$±$\pm 1.2$ $58.8$±$\pm 1.2$ $65.1$±$\pm 1.1$
TAPIR [15] $\mathbf{69.0}$ $\mathbf{66.0}$ $\mathbf{72.0}$
TAPIR [15] $64.1$ $61.3$ $66.9$
HQ-SAM-PT CoTracker [19] $64.8$±$\pm 0.5$ $61.9$±$\pm 0.5$ $67.7$±$\pm 0.5$
Appendix D Different SAM Backbones
The SAM model’s backbone plays an important role in determining its performance and inference speed. In this experiment, we evaluated SAM with different ViT backbones: ViT-Huge (used throughout the work), ViT-Large, and ViT-Base. The results, as measured on the validation subset of DAVIS $2017$ are shown in Tab. 11. Replacing ViT-Huge with ViT-Large results in only a non-significant loss in performance, SAM-PT’s overall performance nevertheless remains non-real-time. Note that the ViT-Huge number in this experiment is slightly different from the one presented in the main manuscript due to the use of different seed values.
Table 11: Performance of SAM-PT with different backbones and their inference speed in semi-supervised VOS on the validation subset of DAVIS $2017$, when using PIPS as the point tracker.
SAM Backbone $\mathcal{J\&F}$ FPS
ViT-Huge $\mathbf{76.7}\pm 0.6$ $1.4$
ViT-Large $76.4\pm 0.6$ $1.8$
ViT-Base $72.2\pm 0.5$ $\mathbf{2.6}$
Appendix E Different SAM Variants
To cater to scenarios where inference speed is crucial, we explored lightweight variants of SAM: Light HQ-SAM [20] and MobileSAM [52]. The performance and speed trade-offs for these variants are summarized in table Tab. 12. Using the HQ-SAM [20] variant of SAM results in the highest performance of $77.64$ points, whereas MobileSAM has the highest inference speed of $5.5$ FPS. Using the lightweight variants doesn’t achieve real-time performance as the bottleneck of the pipeline moves to the point tracker.
Table 12: Performance of lightweight SAM variants and their inference speed in semi-supervised VOS on the validation subset of DAVIS $2017$ when using PIPS [16] as the point tracker.
SAM Variant Backbone $\mathcal{J\&F}$ FPS
HQ-SAM [20] ViT-Huge $\mathbf{77.64}$ $1.3$
SAM [21] ViT-Huge $76.65$ $1.4$
Light HQ-SAM [20] ViT-Tiny $71.30$ $4.8$
MobileSAM [52] ViT-Tiny $71.07$ $\mathbf{5.5}$
Appendix F Detailed Ablation Results
We report detailed experimental results of our ablation studies for configurations using PIPS [16] as the point tracker in Tab. 14 and using CoTracker [19] in Tab. 13.
Table 13: Ablation study on the validation subset of DAVIS $2017$ on the impact of different SAM-PT configurations using CoTracker [19] as the point tracker. PSM: point selection method. PP: positive points per mask. NP: negative points per mask. IRI: iterative refinement iterations. PS: point similarity filtering.
SAM-PT Config. (using CoTracker) DAVIS 2017 Validation [30]
[…]
K-Medoids 16 1 12 $\mathbf{77.6}$±$\pm 0.7$ $\mathbf{74.8}$±$\pm 0.7$ $\mathbf{80.4}$±$\pm 0.7$
Shi-Tomasi 16 1 12 $74.3$±$\pm 0.3$ $72.1$±$\pm 0.3$ $76.5$±$\pm 0.3$
Random 16 1 12 $76.4$±$\pm 1.1$ $73.3$±$\pm 1.1$ $79.4$±$\pm 1.0$
Mixed 16 1 12 $76.4$±$\pm 0.6$ $73.7$±$\pm 0.5$ $79.2$±$\pm 0.6$
K-Medoids 1 1 12 $39.0$±$\pm 0.9$ $36.1$±$\pm 0.9$ $42.0$±$\pm 1.0$
K-Medoids 16 0 12 $76.8$±$\pm 0.7$ $74.1$±$\pm 0.7$ $79.6$±$\pm 0.7$
K-Medoids 16 1 0 $75.3$±$\pm 0.6$ $73.2$±$\pm 0.5$ $77.3$±$\pm 0.6$
K-Medoids 16 1 1 $76.6$±$\pm 0.6$ $74.3$±$\pm 0.6$ $78.9$±$\pm 0.6$
K-Medoids 16 1 100 $77.5$±$\pm 0.7$ $74.7$±$\pm 0.7$ $80.3$±$\pm 0.7$
K-Medoids 16 1 12 $73.8$±$\pm 0.8$ $71.1$±$\pm 0.8$ $76.5$±$\pm 0.8$
Table 14: Detailed ablation study on the validation subset of DAVIS $2017$ on the impact of different SAM-PT configurations using PIPS [16] as the point tracker. PSM: point selection method. PP: positive points per mask. NP: negative points per mask. IRI: iterative refinement iterations. PS: point similarity filtering. RV: reinitialization variant.
SAM-PT Configuration (using PIPS) DAVIS $\mathbf{2017}$ Validation [30]
[…]
(a) point selection method and positive points per mask
Random $1$ 00 00 $37.1$±$\pm 21.7$ $34.3$±$\pm 22.0$ $40.0$±$\pm 21.5$
Random $8$ 00 00 $70.5$±$\pm 1.4$ $68.5$±$\pm 1.4$ $72.6$±$\pm 1.5$
Random $16$ 00 00 $70.0$±$\pm 1.1$ $68.2$±$\pm 1.0$ $71.8$±$\pm 1.2$
Random $72$ 00 00 $62.6$±$\pm 0.4$ $62.3$±$\pm 0.3$ $62.8$±$\pm 0.5$
Shi-Tomasi $1$ 00 00 $20.3$±$\pm 0.1$ $18.3$±$\pm 0.1$ $22.3$±$\pm 0.2$
Shi-Tomasi $8$ 00 00 $72.0$±$\pm 0.3$ $70.3$±$\pm 0.3$ $73.7$±$\pm 0.4$
Shi-Tomasi $16$ 00 00 $66.6$±$\pm 0.4$ $65.7$±$\pm 0.5$ $67.6$±$\pm 0.4$
Shi-Tomasi $72$ 00 00 $54.4$±$\pm 0.3$ $54.8$±$\pm 0.2$ $54.0$±$\pm 0.4$
K-Medoids $1$ 00 00 $32.2$±$\pm 0.7$ $30.4$±$\pm 0.8$ $34.0$±$\pm 0.7$
K-Medoids $8$ 00 00 $\mathbf{72.3}$±$\pm 1.2$ $\mathbf{70.4}$±$\pm 1.3$ $\mathbf{74.3}$±$\pm 1.1$
K-Medoids $16$ 00 00 $71.4$±$\pm 0.2$ $69.8$±$\pm 0.3$ $73.1$±$\pm 0.2$
K-Medoids $72$ 00 00 $58.0$±$\pm 0.2$ $57.3$±$\pm 0.2$ $58.7$±$\pm 0.3$
Mixed $1$ 00 00 $29.9$±$\pm 0.9$ $26.6$±$\pm 0.8$ $33.2$±$\pm 1.4$
Mixed $8$ 00 00 $70.6$±$\pm 0.8$ $68.6$±$\pm 0.8$ $72.5$±$\pm 0.8$
Mixed $16$ 00 00 $\pm 0.7$ $68.2$±$\pm 0.6$ $71.9$±$\pm 0.7$
Mixed $72$ 00 00 $62.8$±$\pm 0.5$ $62.4$±$\pm 0.5$ $63.2$±$\pm 0.6$
(b) negative points per mask
K-Medoids $8$ 00 00 $72.3$±$\pm 1.2$ $70.4$±$\pm 1.3$ $74.3$±$\pm 1.1$
K-Medoids $8$ $1$ 00 $\mathbf{74.1}$±$\pm 0.7$ $\mathbf{72.1}$±$\pm 0.6$ $\mathbf{76.1}$±$\pm 0.7$ $\mathbf{+1.8}$
K-Medoids $8$ $8$ 00 $74.0$±$\pm 0.8$ $71.9$±$\pm 0.8$ $76.0$±$\pm 0.9$
K-Medoids $8$ $16$ 00 $73.4$±$\pm 0.6$ $71.4$±$\pm 0.6$ $75.3$±$\pm 0.6$
K-Medoids $8$ $72$ 00 $72.2$±$\pm 0.4$ $70.3$±$\pm 0.4$ $74.0$±$\pm 0.4$
(c) iterative refinement iterations
K-Medoids $8$ $1$ 00 $74.1$±$\pm 0.7$ $72.1$±$\pm 0.6$ $76.1$±$\pm 0.7$
K-Medoids $8$ $1$ $1$ $75.7$±$\pm 0.7$ $73.4$±$\pm 0.7$ $78.1$±$\pm 0.6$
K-Medoids $8$ $1$ $3$ $76.0$±$\pm 0.6$ $73.4$±$\pm 0.7$ $78.6$±$\pm 0.7$
K-Medoids $8$ $1$ $12$ $\mathbf{76.3}$±$\pm 0.6$ $\mathbf{73.6}$±$\pm 0.6$ $\mathbf{78.9}$±$\pm 0.6$ $\mathbf{+2.2}$
(d) patch similarity filtering
K-Medoids $8$ $1$ $12$ $\mathbf{76.3}$±$\pm 0.6$ $\mathbf{73.6}$±$\pm 0.6$ $\mathbf{78.9}$±$\pm 0.6$ none
K-Medoids $8$ $1$ $12$ $0.002$ $72.7$±$\pm 2.0$ $70.2$±$\pm 1.8$ $75.2$±$\pm 2.1$
K-Medoids $8$ $1$ $12$ $0.01$ $70.7$±$\pm 2.0$ $68.3$±$\pm 1.8$ $73.2$±$\pm 2.1$
(e) point reinitialization
K-Medoids $8$ $1$ 00 A $75.7$±$\pm 0.7$ $73.7$±$\pm 0.6$ $77.7$±$\pm 0.8$
K-Medoids $8$ $1$ 00 B $75.8$±$\pm 0.6$ $73.5$±$\pm 0.9$ $78.1$±$\pm 0.3$
K-Medoids $8$ $1$ 00 C $75.5$±$\pm 0.7$ $73.2$±$\pm 0.8$ $77.8$±$\pm 0.7$
K-Medoids $8$ $1$ 00 D $75.4$±$\pm 0.2$ $73.3$±$\pm 0.2$ $77.5$±$\pm 0.3$
K-Medoids $8$ $1$ $12$ A $76.6$±$\pm 0.8$ $74.0$±$\pm 0.8$ $\mathbf{79.1}$±$\pm 0.8$
K-Medoids $8$ $1$ $12$ B $76.1$±$\pm 0.4$ $73.5$±$\pm 0.5$ $78.6$±$\pm 0.3$
K-Medoids $8$ $1$ $12$ C $75.4$±$\pm 0.6$ $72.8$±$\pm 0.7$ $78.0$±$\pm 0.5$
K-Medoids $8$ $1$ $12$ D $76.4$±$\pm 0.3$ $74.0$±$\pm 0.4$ $78.8$±$\pm 0.3$
K-Medoids $8$ $1$ $12$ $76.3$±$\pm 0.6$ $73.6$±$\pm 0.6$ $78.9$±$\pm 0.6$
K-Medoids $8$ $72$ 00 A $74.9$±$\pm 0.9$ $73.2$±$\pm 0.8$ $76.6$±$\pm 1.0$
K-Medoids $8$ $72$ 00 B $76.0$±$\pm 1.1$ $73.9$±$\pm 1.1$ $78.1$±$\pm 1.1$
K-Medoids $8$ $72$ 00 C $75.1$±$\pm 0.6$ $72.9$±$\pm 0.5$ $77.2$±$\pm 0.7$
K-Medoids $8$ $72$ 00 D $75.6$±$\pm 1.5$ $73.8$±$\pm 1.5$ $77.3$±$\pm 1.6$
K-Medoids $8$ $72$ $12$ A $\mathbf{76.8}$±$\pm 0.7$ $\mathbf{74.5}$±$\pm 0.8$ 79.079.0{79.0}79.0±$\pm 0.6$ $\mathbf{+0.5}$
K-Medoids $8$ $72$ $12$ B $74.8$±$\pm 0.8$ $72.1$±$\pm 0.9$ $77.6$±$\pm 0.7$
K-Medoids $8$ $72$ $12$ C $75.0$±$\pm 0.4$ $72.1$±$\pm 0.4$ $77.8$±$\pm 0.5$
K-Medoids $8$ $72$ $12$ D $75.2$±$\pm 1.1$ $72.7$±$\pm 1.1$ $77.6$±$\pm 1.1$
Appendix G Point Tracking Reinitialization